OpenStudy (anonymous):

M is the midpoint of CF for the points C(2,6) and F(8,10). Find MF Choices: A. 26 B. 2 square root 13 C. square root 13 D. 13 which one ?

OpenStudy (anonymous):

Can you explain what you are trying to find a little better?

OpenStudy (anonymous):

It doesnt have a picture on the question to show. A.\[\sqrt{13}\]

OpenStudy (anonymous):

 Do you think you could help with this one?

OpenStudy (anonymous):

If the distance is what you need than the answer is \[\sqrt{13}\]

OpenStudy (anonymous):

Area of a circle is \[\pi \times r ^{2}\] Therefore, \[A = \pi \times 42^{2}\] A = \[1746\pi \]